OpenStudy (anonymous):

To get home from school, Bob walks four blocks north and three blocks east. What is the straight line distance between Bob’s house and his school?

OpenStudy (anaise):

why can't he just take a bus

TheSmartOne (thesmartone):



TheSmartOne (thesmartone):

So @xav do you know some pythagorean triples? or the pythagorean formula? :)

OpenStudy (anonymous):

ummmm noooo

TheSmartOne (thesmartone):

ok pythagorean theorem: \(\sf a^2 + b^2 = c^2\) c is the hypotenuse for a right triangle so we have 3 and 4 as the legs \(\sf (3)^2 + (4)^2 = (x)^2\) solve for x

OpenStudy (anonymous):

wat

TheSmartOne (thesmartone):

solve for x

OpenStudy (anonymous):

how

TheSmartOne (thesmartone):

what is 3^2 = ?

OpenStudy (anonymous):

9

TheSmartOne (thesmartone):

what is 4^2 = ?

OpenStudy (anonymous):

16

TheSmartOne (thesmartone):

ok what is 9 + 16 = ?

OpenStudy (anonymous):

25

OpenStudy (shadowlegendx):

I would have blown up a pathway with tnt and walked through burning houses. #FastTrips #Creativity

OpenStudy (anonymous):

wtf

TheSmartOne (thesmartone):

lolol xD

TheSmartOne (thesmartone):

ok @xav what is \(\sf\Large\sqrt{25} = ?\)

OpenStudy (anonymous):

idkr

TheSmartOne (thesmartone):

ok what two numbers when multiplied get you 25?

OpenStudy (anonymous):

4 and 3

TheSmartOne (thesmartone):

the same two numbers like \(\sf\Large \sqrt{16} = 4\) because 4 x 4 = 16 another example \(\sf \sqrt{36} = 6\) because 6 x 6 = 36

OpenStudy (anonymous):

i hate math now im droping out

TheSmartOne (thesmartone):

but you just started

OpenStudy (anonymous):

still im droping OUT goodbye MU scolorship

OpenStudy (anonymous):

just can u give me the ansewer

TheSmartOne (thesmartone):

ok tell me if 4 x 4 = 16 and 6 x 6 = 36 and 25 is in between 16 and 36 what number is in between 4 and 6?

OpenStudy (anonymous):

5

TheSmartOne (thesmartone):

and that's your answer

TheSmartOne (thesmartone):

see that wasn't so hard!

OpenStudy (anonymous):

thnkz bro

TheSmartOne (thesmartone):

no need to drop out of math now :D

TheSmartOne (thesmartone):

anytime! :D